OpenStudy (anonymous):

Please help? Will give medal! Determine which system below will produce infinitely many solutions. A; -6x + 3y = 18 4x - 3y = 6 B: 2x + 4y = 24 6x + 12y = 36 C: 3x - y = 14 -9x + 3y = -42 D: 5x + 2y = 13 -x + 4y = -6

OpenStudy (acxbox22):

see which equations are the same

OpenStudy (anonymous):

They're all different...A is x6 B is x=10, y=1 C is x=5, y=1 and C is x=33/11, y=-17/22

OpenStudy (acxbox22):

yes but you can multiply one equation by a number so it becomes another just like a fraction

OpenStudy (anonymous):

A fraction can't be infinitely

OpenStudy (anonymous):

Or are you using that as an example?

OpenStudy (acxbox22):

example

OpenStudy (acxbox22):

for ex. 2x+3y=6 4x+6y=12 these equations are equal because if you multiply the top equation by 2 you get the bottom equation